# Cross-Modality Perturbation Synergy Attack for Person Re-identification

### v2

@@ Abstract
In recent years, there has been significant research focusing on addressing security concerns in single-modal person re-identification (ReID) systems that are based on RGB images. However, the safety of cross-modality scenarios, which are more commonly encountered in practical applications involving images captured by infrared cameras, has not received adequate attention. The main challenge in cross-modality ReID lies in effectively dealing with visual differences between different modalities. For instance, infrared images are typically grayscale, unlike visible images that contain color information. Existing attack methods have primarily focused on the characteristics of the visible image modality, overlooking the features of other modalities and the variations in data distribution among different modalities. This oversight can potentially undermine the effectiveness of these methods in image retrieval across diverse modalities. This study represents the first exploration into the security of cross-modality ReID models and proposes a universal perturbation attack specifically designed for cross-modality ReID. This attack optimizes perturbations by leveraging gradients from diverse modality data, thereby disrupting the discriminator and reinforcing the differences between modalities. We conducted experiments on two widely used cross-modality datasets, namely RegDB and SYSU, which not only demonstrated the effectiveness of our method but also provided insights for future enhancements in the robustness of cross-modality ReID systems.
@@ 1 Introduction
Figure 1: Top 5 predictions of DDAG (one of the cross-modality ReID baseline) before and after our attack on RegDB. Green boxes indicate correctly matched images, while red boxes indicate mismatched images.
@@
With the rapid advancement of surveillance technology, person re-identification (ReID) Yu et al. (2023) has emerged as a pivotal component in the realm of security, garnering escalating attention. ReID constitutes a fundamental task in computer vision, aiming to precisely identify the same individual across diverse locations and time points by analyzing pedestrian images captured through surveillance cameras Zhong et al. (2021). The challenges inherent in this task encompass factors such as changes in viewpoint, lighting conditions, occlusion, and pose variations, culminating in significant appearance variations of the same individual across distinct camera views Zhong et al. (2018).
@@
In traditional ReID, where samples are image-based, the conventional methodology centers on matching visible-to-visible (RGB-to-RGB) data. However, when dealing with diverse scenarios and conditions, especially involving multiple image modalities such as RGB and infrared images, the system needs to intricately handle the differences in images from different modalities. This is essential to ensure that the system exhibits better robustness across different modalities. Hence, cross-modality ReID is considered more challenging due to the need for addressing these modality differences Ye et al. (2022).
Figure 2: The proposed CMPS attack framework. During training, we utilize random grayscale transformations to generate homogeneous grayscale images, bridging the gap between different modalities to assist in learning the universal perturbation. The terms ’optimize’ and ’attack’ denote utilizing gradients from samples of one modality to optimize a universal perturbation, which is then applied to generate adversarial samples on images from another modality for model attacks. By aggregating the gradients of features from different modality images, we cyclically iterate to learn a universal perturbation that can push samples to the common region of the manifold.
1.0.1 Cross-Modality ReID
@@
Cross-modal ReID plays a crucial role in significantly expanding the applicability of traditional ReID methods, focusing on addressing complex matching issues between different image modalities. In practical surveillance systems, the simultaneous use of multiple sensors, such as RGB cameras and infrared cameras, is a common scenario. This task requires innovative solutions to effectively bridge the differences between various modalities, ensuring robust and accurate re-identification of pedestrians in heterogeneous sensor outputs. Taking infrared images as an example, they typically exhibit grayscale features, emphasizing the capture of thermal radiation information, forming a distinct contrast with the rich color information in RGB images. Therefore, when dealing with cross-modal ReID tasks, dual-stream networks are commonly adopted in model design Ye et al. (2022, 2020a).
1.0.2 Security in ReID
Currently, most studies on the security of ReID focus on single-modality systems based on RGB images, and there is insufficient attention to the security of cross-modality ReID systems.
The challenge in cross-modality attacks lies in the substantial visual disparities between different modal inputs, requiring attackers to capture shared features from each modality effectively for perturbation implementation. However, existing attack methods can only generate perturbations by optimizing single-modality information, failing to comprehensively capture shared features among different modalities. Inspired by the design of dual-stream networks in cross-modality ReID, we propose simultaneously utilizing gradients from different modalities to optimize a universal perturbation. This approach allows for capturing shared features among different modalities from a diverse set of samples, enhancing the attack’s universality and applicability to diverse scenarios with modal diversity.
Specifically, we propose the Cross-Modality Perturbation Synergy (CMPS) attack—a novel universal perturbation method tailored for cross-modality ReID. This approach synchronously optimizes perturbations by leveraging gradient information from diverse modality data. In contrast to traditional attack methods that optimize perturbations using information from only one modality, our synergistic approach proves more effective in perturbing cross-modality models. Additionally, during the optimization process, we incorporate homogeneous grayscale transformation to further bridge the gap between different modalities. This aids in synergistically optimizing perturbations, effectively capturing shared features among different modalities. Through this process, it ultimately acquires a universal perturbation capable of pushing each sample into a common region on the manifold Moosavi-Dezfooli et al. (2017), thus better deceiving the discriminator.
@@ 1.0.2 Security in ReID
In our experiments on widely utilized cross-modality ReID datasets, including RegDB Nguyen et al. (2017) and SYSU Wu et al. (2017), we not only showcase the effectiveness of our proposed method but also provide insights for fortifying the robustness of cross-modality ReID systems in the future. This research contributes by bridging gaps in current studies and introducing novel perspectives to study the security challenges in cross-modality ReID systems.
@@
$\bullet$ To the best of our knowledge, our work is the first to investigate vulnerabilities in cross-modality ReID models. Through our method, the cross-modality constraint is explicitly injected into the optimization, improving the universality of the learned cross-modality perturbation.
@@
$\bullet$ Extensive experiments conducted on two widely used cross-modality ReID benchmarks demonstrate the effectiveness of our proposed cross-modality attack. Our method exhibits good transferability even when attacking different models.
@@ 2.0.1 Adversarial Attack
Adversarial attacks are a technique involving the clever design of small input perturbations with the aim of deceiving machine learning models, leading them to produce misleading outputs. This form of attack is not confined to the image domain but extends to models in various fields, including speech and text. Typically, the goal of adversarial attacks is to tweak input data in a way that causes the model to make erroneous predictions when handling these subtly modified samples Goodfellow et al. (2015). In the early stages of research, adversarial attacks had to be customized for each specific sample. However, with the evolution of related studies, universal perturbation attacks were introduced, aiming to find perturbations effective across multiple samples rather than tailored to individual instances Moosavi-Dezfooli et al. (2017). Research on universal perturbation attacks seeks to expose vulnerabilities in models, prompting designers to enhance their robustness to withstand a broader range of adversarial challenges.
2.0.2 Adversarial Attacks in ReID
Some ReID attack methods have been proposed, with current research predominantly focusing on RGB-RGB matching. These methods mainly include:
@@ 2.0.2 Adversarial Attacks in ReID
Metric-FGSM Bai et al. (2020) extends some techniques, inspired by classification attacks, into a category known as metric attacks. These encompass Fast Gradient Sign Method (FGSM) Goodfellow et al. (2015), Iterative FGSM (IFGSM), and Momentum IFGSM (MIFGSM) Dong et al. (2018). The Furthest-Negative Attack (FNA) Bouniot et al. (2020) integrates hard sample mining Hermans et al. (2017) and triple loss to employ pushing and pulling guides. These guides guide image features towards the least similar cluster while moving away from other similar features. Deep Mis-Ranking (DMR) Wang et al. (2020) utilizes a multi-stage network architecture to pyramidally extract features at different levels, aiming to derive general and transferable features for adversarial perturbations. Gong et al. Gong et al. (2022) proposed a method LTA [^2] https://github.com/finger-monkey/LTA_and_joint-defence specifically for attacking color features without requiring additional reference images and discussed effective defense strategies against current ReID attacks. The Opposite-direction Feature Attack (ODFA) Zheng et al. (2023) exploits feature-level adversarial gradients to generate examples that guide features in the opposite direction with an artificial guide.  Yang et al. (2023) introduced a combined attack named Col.+Del. (Color Attack and Delta Attack), which integrates UAP-Retrieval Li et al. (2019) with color space perturbations Laidlaw and Feizi (2019). The inclusion of color space perturbations enhances the attack’s universality and transferability across visible datasets. While this method also explores universal perturbations in ReID, its generality is limited due to the inability to leverage color information in cross-modality problems. In contrast to the aforementioned approaches, our focus lies on addressing cross-modality challenges.
@@ 3 Methodology
In this section, we introduce a universal perturbation designed for cross-modality attacks, referred to as the Cross-Modality Perturbation Synergy (CMPS) attack. Considering the significant differences between different modalities, we introduce a data augmentation method to bridge the gap between modalities, aiding in enhancing the perturbation’s universality across different modalities. Our objective in addressing this problem is to find a universal adversarial perturbation, denoted as $\eta$, capable of misleading the retrieval ranking results of cross-modality ReID models. The adversarial operation involves adding $\eta$ to a query image $I$. The perturbed query image, denoted as $I_{adv}=I+\eta$, is then used to retrieve from the gallery and deceive the cross-modality ReID model $f$. The algorithm is summarized in Alg. 1.
Algorithm 1 Procedure of CMPS attack
Inputs:Visible images $I_{RGB}$ and Infrared (or Thermal) images $I_{ir}$ from dataset $S$, cross-modality ReID model $f$ trained on $S$ , adversarial bound $\epsilon$, momentum value $\theta$, iteration step size $\alpha$, iteration epoch $iter\_epoch$.
Outputs: Cross-modality universal perturbation $\eta$.
1:  Initialize $\delta$ with random noise $\eta$ ( $\|\eta\|_{\infty}\leq\epsilon$), $\Delta^{0}=0$
2:  for $i$ in $iter\_epoch$ do
3:     repeat
4:        $\hat{I_{RGB}}\leftarrow I_{RGB}+\eta$
5:        $\Delta_{RGB}\leftarrow\theta\cdot\Delta^{i-1}+\frac{\partial D(f(\hat{I_{RGB}}),f(I_{RGB}))}{\partial f(\hat{I_{RGB}})}$
6:        $\eta\leftarrow\text{clip}(\eta+\alpha\cdot\text{sign}(\Delta^{i}),-\varepsilon,\varepsilon)$
7:        $\hat{I_{ir}}\leftarrow I_{ir}+\eta$
8:        $\Delta^{i}\leftarrow\theta\cdot\Delta_{RGB}+\frac{\partial D(f(\hat{I_{ir}}),f(I_{ir}))}{\partial f(\hat{I_{ir}})}$
9:        $\eta\leftarrow\text{clip}(\eta+\alpha\cdot\text{sign}(\Delta^{i}),-\varepsilon,\varepsilon)$
10:     until $I_{RGB}$ and $I_{ir}$ are enumerated;
11:  end for
12:  Return $\eta$;
3.0.1 Overall Framework
In Fig. 2, we illustrate the overall framework of the proposed CMPS attack. During the training phase, we suggest optimizing $\eta$ through grayscale augmentation method with images from different modalities to bridge the gap between different modalities, thereby enhancing cross-modality universality. This process retains the structural information of original images while approximating the image style of non-visible modalities. In the attack stage, the acquired $\eta$ deceives reID models, resulting in inaccurate ranking lists. Subsequently, we will provide a detailed introduction to our method.

### v3

@@ Abstract
In recent years, there has been significant research focusing on addressing security concerns in single-modal person re-identification (ReID) systems that are based on RGB images. However, the safety of cross-modality scenarios, which are more commonly encountered in practical applications involving images captured by infrared cameras, has not received adequate attention. The main challenge in cross-modality ReID lies in effectively dealing with visual differences between different modalities. For instance, infrared images are typically grayscale, unlike visible images that contain color information. Existing attack methods have primarily focused on the characteristics of the visible image modality, overlooking the features of other modalities and the variations in data distribution among different modalities. This oversight can potentially undermine the effectiveness of these methods in image retrieval across diverse modalities. This study represents the first exploration into the security of cross-modality ReID models and proposes a universal perturbation attack specifically designed for cross-modality ReID. This attack optimizes perturbations by leveraging gradients from diverse modality data, thereby disrupting the discriminator and reinforcing the differences between modalities. We conducted experiments on three widely used cross-modality datasets, namely RegDB, SYSU, and LLCM. The results not only demonstrate the effectiveness of our method but also provide insights for future improvements in the robustness of cross-modality ReID systems.
footnotetext: Min Jiang and Yunpeng Gong are with the Department of Artificial Intelligence, Key Laboratory of Multimedia Trusted Perception and Efficient Computing, Ministry of Education of China School of Informatics, Key Laboratory of Digital Protection and Intelligent Processing of Fujian and Taiwan Intangible Cultural Heritage, Ministry of Culture and Tourism, Xiamen University, Xiamen 361005, Fujian, P.R. China (e-mail: minjiang@xmu.edu.cn; gongyunpeng@stu.xmu.edu.cn or fmonkey625@gmail.com).
       Corresponding author: Min Jiang
@@ 1 Introduction
With the rapid advancement of surveillance technology, person re-identification (ReID) [38, 50, 9, 24] has emerged as a pivotal component in the realm of security, garnering escalating attention. ReID constitutes a fundamental task in computer vision [18, 35], aiming to precisely identify the same individual across diverse locations and time points by analyzing pedestrian images captured through surveillance cameras [49]. The challenges inherent in this task encompass factors such as changes in viewpoint, lighting conditions [8, 5], occlusion [27, 28], and pose variations, culminating in significant appearance variations of the same individual across distinct camera views [48].
Figure 1: Comparison between traditional and proposed methods: Fig.(a) illustrates traditional attack methods (e.g., FGSM [11], PGD [19]), which are primarily designed for single-modal tasks and lack mechanisms to associate multiple modalities, making them ineffective in simultaneously misleading retrieval results across different modalities. Fig.(b) illustrates the proposed method, which employs an intrinsic mechanism to effectively associate different modalities, thereby misleading retrieval results across multiple modalities simultaneously.
@@
In traditional ReID, where samples are image-based, the conventional methodology centers on matching visible to visible (RGB to RGB) data. However, when dealing with diverse scenarios and conditions, especially involving multiple image modalities such as RGB and infrared images, the system needs to intricately handle the differences in images from different modalities [44, 25, 37]. This is essential to ensure that the system exhibits better robustness across different modalities. Hence, cross-modality ReID is considered more challenging due to the need for addressing these modality differences [41, 42].
@@
Cross-modal ReID [23, 43, 41, 40] plays a crucial role in significantly expanding the applicability of traditional ReID methods, focusing on addressing complex matching issues between different image modalities. In practical surveillance systems, the simultaneous use of multiple sensors, such as RGB cameras and infrared cameras, is a common scenario. This task requires innovative solutions to effectively bridge the differences between various modalities, ensuring robust and accurate re-identification of pedestrians in heterogeneous sensor outputs.
Currently, most research on the security of ReID focuses on single-modality systems based on RGB images [47, 39, 6, 1, 2, 10, 29], while the security of cross-modality ReID systems has received insufficient attention. The challenge in cross-modality attacks arises from significant visual differences among different modality inputs, requiring attackers to effectively capture shared features from each modality for perturbation implementation. However, as shown in Fig. 1, existing attack methods in cross-modal scenarios require optimizing perturbations separately for each modality, lacking an intrinsic mechanism to capture shared knowledge between different modalities, which limits the success rate of the attacks. To address this issue, we propose a synergistic optimization method combined with triplet loss, utilizing information from different modalities to optimize the universal perturbation. This method pushes the features of different samples into a common sub-region that affects the model’s accuracy, as shown in Fig. 2.
Figure 2: Illustration of the CMPS attack framework. We generate homogeneous grayscale images through random grayscale transformations to reduce the differences between modalities, aiding in the learning of a universal perturbation. The process is as follows: first, the gradient from one modality is used to optimize the universal perturbation, which is then applied to another modality’s images to generate adversarial samples for attacks. The new modality’s gradient is then used to further optimize the perturbation and attack the next modality. By aggregating feature gradients from different modalities, we iteratively learn a universal perturbation, pushing samples toward a common region in the manifold. The manifold is represented as a sphere, with identical shapes but different colors representing the same person’s features across modalities. This method captures shared knowledge between modalities, enabling more effective learning of cross-modal universal perturbations.
Specifically, we propose the Cross-Modality Perturbation Synergy (CMPS) method, a universal perturbation approach designed specifically for cross-modality ReID systems. This method simultaneously leverages gradient information from multiple modalities to jointly optimize universal perturbations across visible and infrared images. CMPS incorporates cross-modality triplet loss to ensure feature consistency across different modalities, enhancing the generality of the perturbation. During the synergistic optimization process, CMPS iteratively updates gradients from various modalities within a unified optimization framework, effectively capturing and utilizing shared features across modalities. To further reduce visual differences between modalities, we introduce cross-modality attack augmentation, converting images into grayscale to standardize their visual representation and facilitate the learning of modality-agnostic perturbations. As a result, these universal perturbations push the features of different samples toward a common region in the feature space, significantly diminishing the model’s ability to accurately distinguish identities in cross-modality scenarios, thereby successfully deceiving the model.
@@
In our experiments on widely utilized cross-modality ReID datasets, including RegDB [22], SYSU [34] and LLCM [43], we not only showcase the effectiveness of our proposed method but also provide insights for fortifying the robustness of cross-modality ReID systems in the future. This research contributes by bridging gaps in current studies and introducing novel perspectives to study the security challenges in cross-modality ReID systems.
@@
$\bullet$ To the best of our knowledge, our work is the first to investigate vulnerabilities in cross-modality ReID models. By explicitly incorporating cross-modality constraints into the synergistic optimization process, we enhance the universality of the learned cross-modality perturbations. Additionally, we provide mathematical analysis to demonstrate the superiority of our proposed method over traditional approaches.
@@
$\bullet$ Extensive experiments conducted on three widely used cross-modality ReID benchmarks demonstrate the effectiveness of our proposed cross-modality attack. Our method exhibits good transferability even when attacking different models. The code will be available at https://github.com/finger-monkey/cmps__attack.
In our future work, we will consider incorporating evolutionary search algorithms [32, 14] to capture shared knowledge across different modalities, thereby enhancing the transferability of attacks in any given modality.
@@ 2 Related Works
Adversarial Attack. Adversarial attacks are a technique involving the clever design of small input perturbations with the aim of deceiving machine learning models, leading them to produce misleading outputs. This form of attack is not confined to the image domain but extends to models in various fields, including speech [31] and text [30, 17, 36]. Typically, the goal of adversarial attacks is to tweak input data in a way that causes the model to make erroneous predictions when handling these subtly modified samples [11, 3, 20, 7]. In the early stages of research, adversarial attacks had to be customized for each specific sample. However, with the evolution of related studies, universal perturbation [21] attacks were introduced, aiming to find perturbations effective across multiple samples rather than tailored to individual instances. Research on universal perturbation attacks seeks to expose vulnerabilities in models, prompting designers to enhance their robustness to withstand a broader range of adversarial challenges.
@@
Adversarial Attacks in ReID. Some ReID attack methods have been proposed, with current research predominantly focusing on RGB-RGB matching. These methods mainly include: Metric-FGSM [1] extends some techniques, inspired by classification attacks, into a category known as metric attacks. These encompass Fast Gradient Sign Method (FGSM) [11], Iterative FGSM (IFGSM), and Momentum IFGSM (MIFGSM) [4]. The Furthest-Negative Attack (FNA) [2] integrates hard sample mining [13] and triple loss to employ pushing and pulling guides. These guides guide image features towards the least similar cluster while moving away from other similar features. Deep Mis-Ranking (DMR) [29] utilizes a multi-stage network architecture to pyramidally extract features at different levels, aiming to derive general and transferable features for adversarial perturbations. Gong et al. [6] proposed a method specifically for attacking color features without requiring additional reference images and discussed effective defense strategies against current ReID attacks. The Opposite-direction Feature Attack (ODFA) [47] exploits feature-level adversarial gradients to generate examples that guide features in the opposite direction with an artificial guide. Yang et al. [39] introduced a combined attack named Col.+Del. (Color Attack and Delta Attack), which integrates UAP-Retrieval [16] with color space perturbations [15]. While this method also explores universal perturbations in ReID, its generality is limited due to the inability to leverage color information in cross-modality problems and the lack of a mechanism for associating different modality information. In contrast to the aforementioned approaches, our focus lies on addressing cross-modality challenges.
Algorithm 1 Procedure of CMPS attack
1:Input: Visible images $I_{RGB}$ and infrared (or thermal) images $I_{ir}$ from dataset $S$, cross-modality ReID model $f$ trained on $S$, adversarial bound $\epsilon$, momentum value $\theta$, iteration step size $\alpha$, iteration epoch iter_epoch.
2:Output: Cross-modality universal perturbation $\eta$.
3:Initialize $\eta$ with random noise $\eta\leftarrow Rand(0,1)$, $\Delta^{0}=0$
4:for $i$ in iter_epoch do
5:     repeat
6:         Sample a mini-batch of visible images $I_{RGB}$ and infrared (or thermal) images $I_{ir}$ with $n$ samples
7:         $\hat{I}_{RGB}\leftarrow I_{RGB}+\eta$
8:         Use infrared and grayscale images to compute the triplet loss $L_{RGB}$ for visible images (Eq. 4)
9:         Compute gradient $\Delta_{RGB}$ of $L_{RGB}$ w.r.t. $\eta$:
10:         $\Delta_{RGB}\leftarrow\theta\cdot\Delta^{i-1}+\frac{\partial L_{RGB}}{\partial\eta}$
11:         Update perturbation $\eta$:
12:         $\eta\leftarrow\text{clip}(\eta+\alpha\cdot\text{sign}(\Delta_{RGB}),-\epsilon,\epsilon)$
13:         $\hat{I}_{ir}\leftarrow I_{ir}+\eta$
14:         Use visible and grayscale images to compute the triplet loss $L_{ir}$ for infrared images (Eq. 7)
15:         Compute gradient $\Delta_{ir}$ of $L_{ir}$ w.r.t. $\eta$:
16:         $\Delta^{i}\leftarrow\theta\cdot\Delta_{RGB}+\frac{\partial L_{ir}}{\partial\eta}$
17:         Update perturbation $\eta$:
18:         $\eta\leftarrow\text{clip}(\eta+\alpha\cdot\text{sign}(\Delta^{i}),-\epsilon,\epsilon)$
19:     until all mini-batches are processed
20:end for
21:return $\eta$
@@ 3 Methodology
In this section, we introduce a universal perturbation designed for cross-modality attacks, referred to as the Cross-Modality Perturbation Synergy (CMPS) attack. Considering the significant differences between different modalities, we propose a attack augmentation method to bridge the gap between modalities, aiding in enhancing the perturbation’s universality across different modalities. Our objective in addressing this problem is to find a universal adversarial perturbation, denoted as $\eta$, capable of misleading the retrieval ranking results of cross-modality ReID models. The adversarial operation involves adding $\eta$ to a query image $I$. The perturbed query image, denoted as $I_{adv}=I+\eta$, is then used to retrieve from the gallery and deceive the cross-modality ReID model $f$. The algorithm is summarized in Alg. 1.
3.1 Overall Framework
In Fig. 2, we illustrate the overall framework of the proposed CMPS attack. During the training phase, we optimize $\eta$ using our cross-modality attack augmentation method, which leverages images from different modalities to bridge their inherent differences and enhance cross-modality universality. In the attack phase, the optimized $\eta$ deceives reID models, leading to inaccurate ranking lists. Section 3.2 outlines the framework and overall optimization objective, providing a macro-level overview. Section 3.4 delves into the specific process of perturbation optimization across different modalities.
3.2 Optimizing Loss Functions for Attacking
Our study aims to deceive cross-modality ReID models using a universal perturbation. We have specifically designed a triplet loss tailored for our proposed attack method, which can correlate different modalities and influence the distance relationships between images from different modalities.
We follow the approach of [16] to optimize the perturbation using cluster centroids. This method directly impacts the similarity between pedestrian identities in the ReID model’s feature space (rather than the similarity between individual samples), making it more effective. Subsequently, leveraging the acquired cluster centroids, we apply our triplet loss to distort the pairwise relations between pedestrian identities. This process can be represented as follows:
[…]
As shown in Fig. 4, the loss function mentioned above fully leverages the triplet-wise relationships across different modality. Through this loss, we are able to pull the negative samples of each modality closer to the adversarial samples and push the positive samples of each modality away from the adversarial samples. Here, $C^{p}_{v}$ and $C^{n}_{v}$ represent the cluster centroids of the positive samples to push and negative samples to pull, respectively, in the original visible (RGB) image feature space of the training data. Similar definitions apply to other modalities. $f^{adv}_{v}$, $f^{adv}_{g}$, and $f^{adv}_{ir}$ denote the perturbed features of the disturbed image in the visible, grayscale, and infrared (or thermal) modalities, respectively.
3.3 Cross-Modality Attack Augmentation Method
Intuitively, as illustrated in Fig. 4, maximizing the overlap of common factors across different modalities facilitates the capture of shared features by the learned perturbation. Grayscale images, being inherently homogeneous, serve as effective mediators between diverse modalities. Consequently, we introduce random grayscale transformations into adversarial attack methods, referred to as Cross-Modality Attack Augmentation. This approach guides cross-modality perturbations by leveraging homogeneous grayscale images sourced from diverse modalities. The primary objective is to explore the underlying structural relationships across heterogeneous modalities.